OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth.

OpenStudy (amistre64):

how long will it take to double? ln(2)/r

OpenStudy (amistre64):

A = P e^rt A/P = e^rt ln(A/P) = rt ln(A/P) ----- = t r

OpenStudy (anonymous):

yes i want to round the answer to the nearest tenth

OpenStudy (amistre64):

once you plug that into a calculator, just rnd it off them